The sum 1+3+5+7+...........+29+30+31+32+..........+60=
 
 
Please explain 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
we haveve formula of sum of n natural numbers from starting
 
sigma n = n (n +1 ) /2
 
hence put n = 60
 
 = 60 * 61/2 = 30* 61 = 1830